# The Physics of Preference: Unravelling Imprecision of Human Preferences through Magnetisation Dynamics

### v1

@@ Abstract
Paradoxical decision-making behaviours such as preference reversal often arise from imprecise or noisy human preferences. By harnessing the physical principle of magnetisation reversal in ferromagnetic nanostructures driven by electric current, we developed a model that closely reflects human decision-making dynamics. Tested against a spectrum of psychological data, our model adeptly captures the complexities inherent in individual choices. This blend of physics and psychology paves the way for fresh perspectives on understanding human decision-making processes.
@@
Introduction—Theoretical and experimental research in economics and psychology posits that human preferences are inherently noisy and imprecise [1] (see [2, 3] for a definition of noise and imprecision). For example, such phenomena as Allais Paradox, Ellberg Paradox, Endowment Effect and Preference Reversal (PR) demonstrate that human decisions deviate from predictions of traditional rational models [4]. These deviations follow certain patterns [3]. While the classical theory suggests that individuals consistently maximise their utility based on stable preferences, experiments present situations where choices are inconsistent or even contradictory [5, 6].
@@
PR was first observed in the gambling studies involving [7]: (i) a choice task, where people were asked to decide between a relatively safe lottery with a low payoff (Probability- or P-bet) and a relatively risky lottery with a high payoff (Dollar- or \$-bet) as well as (ii) a valuation task, where the same people were asked to indicate lottery prices if they were to sell P-bet and \$-bet. The observed people’s preferences were often contradictory: opting for the P-bet they put a higher price on the \$-bet.
Alongside other behavioural irregularities, PR has reshaped contemporary discourse in both economics and psychology. Traditional models such as Expected Utility Theory (EUT), which held that individual choices were reflective of consistent and stable preferences, have been challenged by empirical findings [8]. The existence of such paradoxes implied that human decision-making was more multifaceted than previously assumed.
Understanding and accurately predicting these anomalies is not merely an academic exercise. As policy-makers strive to devise strategies and interventions that depend on opinion polls, the need for models predicting human behaviour is pressing. Yet, capturing the essence of phenomena like PR within traditional models remains a challenge. For example, transitivity posits that if an individual prefers $A$ to $B$ and $B$ to $C$, then they should prefer $A$ to $C$. However, PR defies the transitivity, suggesting that human preferences might be imprecise.
Figure 1: (a) Idealised deterministic (the dotted line) and probabilistic (the line with circular markers) PR curves. (b) Simulated magnetisation reversal behaviour for different values of the static magnetic field $H_{0}$.
This instigates debates on the need to (i) reconsider some axioms of decision theory [9]; (ii) devise theories relaxing transitivity assumption [10]; (iii) embed decision theories into models of noise-synthesising stochastic versions of known theories [5]; and (iv) develop models that assume that preferences are inherently imprecise [3].
@@
The latter approach has demonstrated that individual choices under risk are probabilistic rather than deterministic, i.e. the same person may make different choices when presented with the same problem multiple times. Experiments show that people’s decisions to accept or refuse sure payoff amounts in exchange for a risky lottery varies from consistently selecting bet over cash (when the sure payoff amounts offered against the lottery are not sufficiently attractive) to consistently selecting cash over bet (when the sure payoff amounts increase) with some imprecision interval in-between, when individuals make probabilistic choices between cash and the bet [11, 12, 3].
@@
While these debates are ongoing, it has been noticed that physical processes can adequately describe many psychological effects and decision-making phenomena. For example, there exist classical physical models of confirmation bias [13], public opinion formation [14, 15, 16] and backfire effect [17]. Quantum-mechanical models of decision-making [18, 19, 20] that originate from the quantum mind hypothesis [21, 22] have also been developed.
However, there remains an imperative to juxtapose the predictions of such models against empirical findings. In this pursuit, our paper not only elucidates the connections between physics and decision theory, but also provides a comprehensive evaluation of a real-life physical model against rich empirical datasets. Through this endeavour, we aim to herald a paradigm shift in understanding and predicting human decision-making.
@@
Study 1—In a study of imprecision intervals [11, 12, 3], an individual is given a number of choices between a lottery $B$ and a series of increasing sure payoff amounts $A_{j}$$j=0,\dots,13$. Since each choice is made multiple times and independently of the previous ones, we model this individual’s preferences as a probability distribution.
@@
Figure 1a shows two theoretically possible distributions. The dotted line denotes the deterministic preferences of an individual who chooses $B$ when $j<6$, never chooses $B$ when $j>6$ and is exactly indifferent between $A$ and $B$ when $j=6$. The curve with the circular markers denotes the probabilistic preferences. When the sure thing $ST\leq A_{3}$, the individual always chooses $B$. When $ST\geq A_{10}$, the individual never chooses $B$. For $A_{3}<A<A_{10}$ there is an imprecision interval.
@@
The probabilistic distribution in Fig. 1a is drawn as a sigmoid-like curve since this function robustly fits many intuitions [3] and datasets [11, 12, 3, 5, 9]. Figure 2 shows five examples of experimental curves (median patterns) plotted using the data obtained from the laboratory decision-making experiments with human subjects.
@@
All experiments were conducted in the University of Warwick using subject pools of undergraduate, postgraduate and executive students over the age of 18. In Fig. 2a, 81 subjects were making a series of binary choices between a lottery which gave £40 with probability 80% or 0 with probability 20% and sure payoff amounts between £16 and £36 with a step of £2. In Fig. 2b, 101  subjects were offered a lottery which yielded £40 with probability 30% or 0 with probability 70% versus sure amounts of money between £4 and £12 with a step of £1. In Fig. 2c, 101 experimental participants make choices between a lottery providing £15 with probability 70% and 0 otherwise against sure amounts between £4 and £12 with a step of £1. In Fig. 2d, 184 subjects face a choice between a chance of winning £12 with probability 80% and sure amounts from £3 to £11 with a step of £1. In Fig. 2e, 184 people decide between £15 with probability 25% and sure monetary amounts between £3 to £11 with a step of £1. In all lotteries, each person made each of the binary choices 4 times. We can see that all five examples demonstrate robust evidence in favour of the existence of an imprecision interval, where people make probabilistic choices between a lottery and sure monetary amounts.
Figure 2: Experimental PR data of the five lotteries. The straight dotted lines are the prediction of EUT under the assumption that the participants are risk-neutral.
@@
To model the experimental decision-making behaviour, we employ magnetisation dynamics in ferromagnetic metal (FM) structures. In general, an electric current is unpolarised since it involves electrons with a random polarisation of spins. However, a current that passes through a thin FM film with a fixed magnetisation direction can become spin-polarised since spins become oriented predominantly in the same direction [23]. This physical effect is exploited in a spin transfer torque nano-oscillator (STNO), where a layer with a fixed direction of magnetisation ${\bf M}$ is separated from a thinner FM layer by a non-magnetic metal layer [24]. When a spin-polarised current is directed from the “fixed” magnetisation layer into the “free” magnetisation layer, the static equilibrium orientation of magnetisation in the “free” layer becomes destabilised. Depending on the strength of the electric current, the destabilisation of magnetisation can lead to either stable precession of magnetisation of the “free” layer about the direction of the effective magnetic field or to a reversal of the magnetisation direction (Fig. 3).

### v2

@@ Abstract
Paradoxical decision-making behaviours such as preference reversal often arise from imprecise or noisy human preferences. Harnessing the physical principle of magnetisation reversal in ferromagnetic nanostructures, we developed a model that closely reflects human decision-making dynamics. Tested against a spectrum of psychological data, our model adeptly captures the complexities inherent in individual choices. This blend of physics and psychology paves the way for fresh perspectives on understanding the imprecision of human decision-making processes, extending the reach of the current classical and quantum physical models of human behaviour and decision-making.
keywords:
Preference reversal, Magnetisation reversal, Quantum mind, Decision-making
[…]
[label1]organization=Artificial Intelligence and Cyber Futures Institute, Charles Sturt University,city=Bathurst, postcode=2795, state=NSW, country=Australia
[…]
[label2]organization=The Alan Turing Institute, British Library,city=London, postcode=NW1 2DB, country=United Kingdom
[…]
[label3]organization=The University of Sydney Business School,city=Darlington, postcode=2006, state=NSW, country=Australia
1 Introduction
@@
Theoretical and experimental research in economics and psychology posits that human preferences are inherently noisy and imprecise Lin71 (see Bha17 ; loomes2017preference ; blavatskyy2021probabilistic ; busemeyer2019primer ; sharma2023scarcity for definitions of noise and imprecision as well as detailed reviews of the relevant literature). For example, such phenomena as Allais Paradox, Ellberg Paradox, Endowment Effect and Preference Reversal (PR) demonstrate that human decisions deviate from predictions of traditional rational models moffatt2009experimental . These deviations follow certain patterns loomes2017preference . While the classical theory suggests that individuals consistently maximise their utility based on stable preferences, experiments present robust evidence that human choices under risk and uncertainty are often inconsistent or even contradictory chater2018mind ; blavatskyy2010models ; bardsley2009experimental .
@@
PR was first observed in the gambling studies involving Lic71 : (i) a choice task, where people were asked to decide between a relatively safe lottery with a low payoff (Probability- or P-bet) and a relatively risky lottery with a high payoff (Dollar- or \$-bet) as well as (ii) a valuation task, where the same people were asked to indicate lottery prices if they were to sell P-bet and \$-bet. The observed people’s preferences were often paradoxical: while opting for the P-bet in the binary choice, they were putting a higher price on the \$-bet, thereby indicating a preference for P-bet in a simple two-option choice, but preference for \$-bet when stating their certainty equivalents for both options.
@@
Alongside other behavioural irregularities (see, e.g., bardsley2009experimental ; blavatskyy2023common ), PR has reshaped contemporary discourse in both economics and psychology. Traditional deterministic models such as Expected Utility Theory (EUT), which held that individual choices were reflective of consistent and stable preferences, have been challenged by empirical findings Gre79 ; loomes2017preference . Furthermore, recent research demonstrated that preference reversals constitute a challenge for many stochastic decision theories (e.g., when decision theories such as Expected Utility Theory or Cumulative Prospect Theory are embedded into models of stochastic choice such as Fechner model) butler2007imprecision ; loomes2017preference . The latter approach has demonstrated that individual choices under risk are probabilistic rather than deterministic, i.e. the same person may make different choices when presented with the same problem multiple times. Experiments show that people’s decisions to accept or reject sure payoff amounts in exchange for a risky lottery varies from consistently selecting bet over cash (when the sure payoff amounts offered against the lottery are not sufficiently attractive) to consistently selecting cash over bet (when the sure payoff amounts increase) with some imprecision interval in-between, when individuals make probabilistic choices between cash and the bet mosteller1951experimental ; butler2007imprecision ; loomes2014testing ; loomes2017preference . The existence of such paradoxes implied that human decision-making was more multifaceted than previously assumed.
Understanding and accurately predicting these anomalies is not merely an academic exercise. As policy-makers strive to devise strategies and interventions that depend on opinion polls, the need for models predicting human behaviour is pressing. Yet, capturing the essence of phenomena like PR within traditional deterministic and stochastic models remains a challenge. For example, many deterministic and stochastic decision theories honour the transitivity axiom, which posits that if an individual prefers $A$ to $B$ and $B$ to $C$, then they should prefer $A$ to $C$. However, PR defies the transitivity, suggesting that human preferences might be imprecise butler2007imprecision ; butler2018predictably .
Figure 1: (a) Idealised deterministic (the dashed line) and probabilistic (the line with circular markers) PR curves. Examples of real-life data that resemble the shape of the idealised curve are given in Fig. 4. (b) Representative simulated magnetisation reversal behaviour for different values of the static magnetic field $H_{0}$. Additional physical details are presented in the main text and the caption to Fig. 5. The shape of the probabilistic curve in panel (a) closely resembles the magnetisation curves in panel (b). Either a single magnetisation reversal curve or a judicious combination of several magnetisation reversal curves, as demonstrated in the main text, can approximate both experimental curves in Fig. 4 and more complex dataset in Fig. 6.
This instigates debates on the need to (i) reconsider some axioms of decision theory butler2018predictably ; blavatskyy2021probabilistic ; (ii) devise theories relaxing transitivity assumption loomes1982regret ; (iii) embed decision theories into models of noise-synthesising stochastic versions of known theories blavatskyy2010models ; and (iv) develop further models that assume inherently imprecise human preferences loomes2017preference .
Figure 2: Location of the proposed physical Magnetisation Reversal model on a diagram that evaluates the existing decision theory models using a 1–5 scale applied to such model characteristics as Predictive Capability, Applicability, Complexity and Interpretation, Preference Reversal Explanation and Empirical Validation. To obtain the relative scoring, we first compiled a detailed table of decision theories based on the comprehensive literature review and the content of the relevant papers. We then assigned a score based on table entries and plotted the resulting diagram based on the obtained scores. This diagram does not reflect the fact that the Magnetisation Reversal model is physically phenomenological, was developed independently of any other decision theory model and, de facto, is unique as further demonstrated in Fig. 3. The software and raw data used to prepare Fig. 2 can be accessed following the instructions given in the Data Availability section.
@@
While these debates are ongoing, it has been noticed that physical processes can adequately describe many psychological effects and decision-making phenomena (e.g., quantum logic has been previously applied by social psychologists to model decision making under risk and uncertainty using models of quantum cognition yearsley2016quantum ; busemeyer2019primer ). Furthermore, there exist classical physical models of confirmation bias Gro17 , public opinion formation Gal05 ; Cas09 ; Red19 , preference Cap21 ; Kva21 and backfire effect Hoh23 . Thermodynamic models of decision-making where information processing is represented as changes in states of a thermodynamical system and quantified by differences in free energy have also been proposed Ort13 ; Pak13 ; Eva21 ; Ann22 .
Quantum-mechanical models of decision-making Atm10 ; Aer14 ; Bus12 ; Aln17 that originate from the quantum mind hypothesis Khr06 ; mindell2012quantum have been developed following a suggestion that the human mental states can be represented as a combination of two binary states ‘0’ and ‘1’ Atm10 ; Bus12 ; Pot22 . Mathematically and physically, such a representation corresponds to a superposition of the basis quantum-mechanical states $|0\rangle$ and $|1\rangle$ employed in quantum computers.
Instead of binary bits, a quantum computer uses a quantum bit (qubit) that corresponds to the states $|0\rangle=\left[\begin{smallmatrix}1\\ 0\end{smallmatrix}\right]$ and $|1\rangle=\left[\begin{smallmatrix}0\\ 1\end{smallmatrix}\right]$. Unlike the ‘0’ and ‘1’ binary states of a classical digital computer, the states of a qubit are a superposition $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ with $|\alpha|^{2}+|\beta|^{2}=1$. As part of a quantum measurement, a qubit interacts in a controlled way with an external system from which the state of the qubit under measurement can be obtained. For instance, using the projective measurement operators $M_{0}=|0\rangle\langle 0|$ and $M_{1}=|1\rangle\langle 1|$ Nie02 , the measurement probabilities for $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ become $P_{|0\rangle}=|\alpha|^{2}$ and $P_{|1\rangle}=|\beta|^{2}$. This measurement outcome means that the qubit will be in one of its basis states.
Quantum computational algorithms are exponentially faster than any possible classical computational algorithm Nie02 . Therefore, it has been shown that quantum mechanics can explain certain psychological and decision-making processes better than any classical model Pot09 ; Bus12 ; Pot22 . Those theoretical findings have been convincingly supported by a large and growing body of experimental research works that have examined diverse paradoxical phenomena of human behaviour such as Allais Paradox and Ellsberg Paradox Aer11 ; Aer12 ; Blu16 ; Wei19 ; Ish20 ; Ost22 ; Abe23 .
In this paper, we further extend the ability of physical models of decision-making to capture the effect of preference imprecision which often results in the preference reversals. In this pursuit, our paper not only makes a completely new connection between physics and decision theory, which has not been explored in the previous literature, but also provides a comprehensive evaluation of a real-life physical model against rich empirical datasets. In particular, apart from several traditionally designed decision-theoretic laboratory experiments involving lotteries, we validate our model using a dataset obtained from the “Deal or No Deal” video game based on the famous TV show.
As shown in Fig. 2, while our Magnetisation Reversal model is capable of explaining phenomena previously discussed in decision-theoretic literature, it makes a novel contribution and is based on completely different core principle contrasting to other theories. Given the novelty of our approach, the whole Fig. 2 should be viewed mostly as an attempt to find a tentative place for our model in a general taxonomy of the existing decision theory models. For instance, although in Fig. 2 we place our model near the quantum models of decision-making, strictly speaking our model is classical since the main equation of our model—the Landau-Lifshitz-Gilbert equation—is a classical phenomenological model that describes the quantum-mechanical phenomenon of magnetisation Lak11 .
By definition, a phenomenological model seeks to represent the essential behaviour observed in the experiment not necessarily attempting to elucidate the origin of the effect using the existing theories Mcm68 . Hence, our phenomenological model is concerned with reproducing certain decision-making data as they appear in our experience (e.g., visually) using the physical effect of magnetisation reversal. Specifically, we choose this effect because the magnetisation reversal curves closely resemble empirical decision-making data discussed in this paper (e.g., compare the shapes of the probabilistic curve in Fig. 1a and the magnetisation curves in Fig. 1b).
Thus, we leave it up to the readers to make either a further connection between our model and the other models or classify our model as a new class of models of decision-making. To help the readers in navigating the literature, in Fig. 3 we visualise the links of our work to the previous relevant publications. We can see that our theoretical and computational approach stands out from the plethora of competing models reported in the literature.
Figure 3: Literature analysis-based taxonomy of decision making models complementing the data presented in Fig. 2. The connection chart confirms the core-principle independence of our model (shown in red) of the previous theoretical approaches and establishes our model’s distinctiveness compared to other models. The graph demonstrates that while this paper deals with the same behavioural concepts as the previous literature (preference reversals), our approach is different from previously proposed in the literature. The network graph was produced in Python programming language by constructing a similarity index between the papers cited in this manuscript, based on keywords. To ensure robustness, results of our analysis for a subset of papers were cross-checked using the online Connected Papers tool.
2 Study 1
2.1 Empirical Data and Experimental Procedures
@@
In several relevant studies recording imprecision intervals, mosteller1951experimental ; butler2007imprecision ; loomes2017preference , an individual is given a number of choices between a lottery $B$ and a series of increasing sure payoff amounts $A_{j}$$j=0,\dots,13$. Since each choice is made multiple times and independently of the previous ones, we model this individual’s preferences as a probability distribution.
@@
Figure 1a shows two theoretically possible distributions. The dashed line denotes the deterministic preferences of an individual who chooses $B$ when $j<6$, never chooses $B$ when $j>6$ and is exactly indifferent between $A$ and $B$ when $j=6$. The curve with the circular markers denotes the probabilistic preferences. When the sure thing $ST\leq A_{3}$, the individual always chooses $B$. When $ST\geq A_{10}$, the individual never chooses $B$. For $A_{3}<A<A_{10}$ there is an imprecision interval.
@@
The probabilistic distribution in Fig. 1a is drawn as a sigmoid-like curve since this function robustly fits many intuitions loomes2017preference and datasets mosteller1951experimental ; loomes2014testing ; loomes2017preference ; blavatskyy2010models ; butler2018predictably . To provide a specific example of data resulting in sigmoid-like functions, establishing the robustness of the observed behavioural patterns (sigmoid-shaped curves) in Fig. 4 we present five examples of experimental curves plotted using the data obtained from the laboratory decision-making experiments with human subjects. We plot the median patterns in the left column of Fig. 4 and present their respective cumulative distributions in the right column. The Cumulative Distribution Function (CDF) represents the probability that a subject’s propensity to stick with a lottery is less than or equal to a certain level as the monetary stakes change as described below.
All experiments were conducted in the University of Warwick using subject pools of undergraduate, postgraduate and executive students over the age of 18 (no minors participated in the experiments). Upon arrival to the laboratory, each participant was provided with the study information sheet and a consent form. Informed consent was obtained from all participants in the study.
@@
In Fig. 4a, 81 subjects were making a series of binary choices between a lottery which gave £40 with probability 80% or 0 with probability 20% and sure payoff amounts between £16 and £36 with a step of £2. In Fig. 4b, 101  subjects were offered a lottery which yielded £40 with probability 30% or 0 with probability 70% versus sure amounts of money between £4 and £12 with a step of £1. In Fig. 4c, 101 experimental participants make choices between a lottery providing £15 with probability 70% and 0 otherwise against sure amounts between £4 and £12 with a step of £1. In Fig. 4d, 184 subjects face a choice between a chance of winning £12 with probability 80% and sure amounts from £3 to £11 with a step of £1. In Fig. 4e, 184 people decide between £15 with probability 25% and sure monetary amounts between £3 to £11 with a step of £1. In all lotteries, each person made each of the binary choices four times. We can see that all five examples demonstrate robust evidence in favour of the existence of an imprecision interval, where people make probabilistic choices between a lottery and sure monetary amounts.
Thus, Fig. 4 plots summary patterns for 81, 101, 101, 184 and 184 experimental participants, respectively. Considering that in the panels (a)–(e) of Fig. 4 each participant answered at least nine binary choice questions that were repeated four times (hence, we have at least 36 points per participant per case), we additionally plotted the response curves as well as the utility functions for each individual separately, also calculating the number of individuals in each of the cases (a)–(e), who demonstrated patterns consistent with having an imprecision interval according to our definition. Overall, our calculations reveal that in each case the majority of participants exhibited patterns consistent with having an imprecision interval. Importantly, 73% (59 out of 81), 73% (74 out of 101), 90% (91 out of 101), 98% (181 out of 184) and 93% (172 out of 184) of participants had patterns consistent with imprecision intervals in the cases labelled as (a), (b), (c), (d) and (e), correspondingly. Raw data relevant to Fig. 4 can be accessed following the instruction given in the Data Availability section.
Figure 4: Experimental PR data of the five lotteries (each row of the figure corresponds to an independent dataset). The left column (panels (a.i)–(e.i)) shows the median patterns, but the right column (panels (a.ii)–(e.ii)) presents the respective cumulative distributions. Cumulative probability is abbreviated as ’Cumul. Prob.’, CDF stands for cumulative distribution function. The straight dashed lines are the prediction of EUT under the assumption that the participants are risk-neutral. The EUT predictions are provided for illustration only and are not used to validate the magnetisation reversal model. In each row, we first (a.i–e.i) demonstrate the median propensity to stick to a lottery against a series of sure monetary amounts displayed on the horizontal axis (to that end, the number of choices to stick with the lottery was summed across four repetitions for the same subject and median probability to stick with the lottery was calculated across all subjects). On the right hand side (a.ii–e.ii), we demonstrate the corresponding cumulative distribution functions reflective of the mean propensity to stick with the lottery against a series of stakes.
2.2 Magnetisation Reversal Model
@@
To simulate the experimental decision-making behaviour, we employ magnetisation dynamics in ferromagnetic metal (FM) structures. In general, an electric current is unpolarised since it involves electrons with a random polarisation of spins. However, a current that passes through a thin FM film with a fixed magnetisation direction can become spin-polarised since spins become oriented predominantly in the same direction Ral08 . This physical effect is exploited in a spin transfer torque nano-oscillator (STNO), where a layer with a fixed direction of magnetisation ${\bf M}$ is separated from a thinner FM layer by a non-magnetic metal layer Rip04 . When a spin-polarised current is directed from the “fixed” magnetisation layer into the “free” magnetisation layer, the static equilibrium orientation of magnetisation in the “free” layer becomes destabilised. Depending on the strength of the electric current, the destabilisation of magnetisation can lead to either stable precession of magnetisation of the “free” layer about the direction of the effective magnetic field or to a reversal of the magnetisation direction (Fig. 5).
We model the dynamics of magnetisation using the Landau-Lifshitz-Gilbert equation Sla09 :
where $\gamma$ is the gyromagnetic ratio. The first term of the right-hand side of Eq. (1) governs the precession of $\bf M$ (Fig. 5a) about the direction of the effective magnetic field $\textbf{H}_{eff}=H_{0}\textbf{e}_{\rm{z}}+\textbf{H}+\textbf{H}_{ex}$, where $H_{0}$ is the external static magnetic field orientated along the z-axis, H is the dynamic field due to currents and magnetic sources such as demagnetising field and eddy current fields and $\textbf{H}_{ex}$ is the exchange field Mak13 . The dissipative torque is Sla09
where $M_{s}$ is the saturation magnetisation of the “free” magnetisation layer and $\alpha_{G}$ is Gilbert damping parameter Mak13 ; Sla09 . The Slonczewski-Berger torque is
where $I$ is the strength and $\hat{e}_{p}$ is the direction of the spin polarisation of the current. The parameter $\sigma_{0}$ incorporates fundamental physical constants and the thickness of the “free” magnetisation layer Sla09 .
Equation (1) is solved consistently with Maxwell’s equations using a finite-difference time-domain (FDTD) method and material parameters presented in Mak13 . We model a point-contact STNO structure experimentally studied in Rip04 . Since we employ a one-dimensional FDTD scheme, the strength of the current $I$ required to produce a magnetisation reversal in our model is lower than that in Rip04 .